A clarinet player hears the warbling of beats twic
Answer #1

1. apply beat frequency f = fa - fb

given fbeat = 2

fb = 440 Hz

so fa = 440 + 2   = 442 Hz

-------------------------------

2. apply the law of conservation of momentum as I1 W1 = I2 W2

I is moment of inertia   = mr^2

I1 = I

i2 = I/4

but W remains constant

so

option A is true